OpenStudy (aliaaalihassan):

x^3+2x+4=12 How can I factorize this to give me X=2 ?

OpenStudy (cwrw238):

f(x) = x^3 + 2x - 8 = 0 if x = 2 is a root of this equation then f(2) = 0 but f(2) = 8+4 - 8 = 4 so 2 is not a root

OpenStudy (cwrw238):

r u sure the first term is x^3? - because if it's x^2 then x=2 is a root

OpenStudy (aliaaalihassan):

sorry I meant x^3 +2x -12 = 0

OpenStudy (cwrw238):

ok then f(2) = 0 which means that x-2 is a factor you can use long division:

OpenStudy (cwrw238):